OpenStudy (anonymous):

Last year you mowed grass or shoveled snow for 12 households. You charged $200 a season to mow a yard and $180 a season to shovel snow. You earned $2260 last year. How many households did you mow?

OpenStudy (anonymous):

Let s= yards snow shoveled; Let m= yards mowed. For two unknowns, you need two equations to solve. So s+m=12 $200m+$180s= $2,260 Reagrange the first equation. So s=12-m Plug this equation into the second one. $200m+($180)(12-m)=$2,260 $200m-$180m+ $2,160=$2,260 $20m=$2,260-$2,160 m=5 So he mowed 5 lawns.